OpenStudy (anonymous):

√2x+3 (all rooted) -7=0

OpenStudy (anonymous):

√2x+3 is all rooted not -7=0

OpenStudy (dumbcow):

\[\sqrt{2x+3} = 7\] \[2x+3 = 49\] \[x = 23\]